Crew 235 – Final Mission Summary

MDRS Crew 235 Spaceward Bound Mission Summary

May 7, 2021

Summary Title: Generally Speaking, this is how it went

Authors’ names: Thomas Quayle (Science Lion Simba) and Jen Carver-Hunter

Sol Activity Summary:

At the start of our mission, nine strangers came together with different hopes and expectations, all with the same end goal – come together with other like-minded individuals within the education community and emerge with new skills or ideas. Amazingly, every step of our journey together built the outcome that would meet and exceed everyone’s expectations.

Many of the crew interactions and experiences have been recorded in the journal reports that have been submitted each day of the mission. This report will focus on the accomplishment of our mission goals, which included completing a microbiology experiment and geologic fieldwork for collection, comparison, and analysis.

For our microbiology experiment, crew prepared agar, which they used to grow the fungi and bacteria that were swabbed from different areas inside the Hab. After allowing the specimens to incubate for 3 days, crew sorted and categorized the agar to identify the locations that were swabbed. This was done in a blind categorization so both crew teams approached it through flexible problem solving. The first team used a matrix system for identification while the second team used process of elimination and benchmarks to identify the tested locations for each sample. Every crew member felt the microbiology experiment was a success in that it would be transferable to a K-12 classroom environment.

For our geologic fieldwork experience, crew members used a presentation about Mars geology by Dr Rupert as a spring board for identification of MDRS rock samples that represent the geology of Mars. We then used 3 EVAs over the next 3 days to collect rock samples we will use to create our MDRS-Mars rock collections. On the first EVAs, crew members traveled to Kissing Camels Ridge and followed the ridge to the west to collect petrified wood, which we were using as an analogous sample for the sulphur located on Mars. Crew members were delighted to find a diverse variety of samples, including a wide range of colors and textures. On the second EVAs, crew members traveled to Barainca Butte to collect lava, basalt, and granite samples. We were also able to collect gryphaea samples during this EVA as an added bonus! The second stop for this EVA was near Zubrin’s head to collect sandstone blueberries from a blueberry nursery. This geologic phenomenon was one that continued to delight and amaze crew members. On the third set of EVAs, crew members traveled to Copernicus Overlook to see the big picture of MDRS geology. After the overlook, crews returned east on Brahe Hwy, looking for the glint of gypsum near Beige Moon. Every crew member was surprised and excited to see the glinting shine that covered the ground since it was not observable on the way to the overlook. Once gypsum was collected, crews continued on to Cowboy Corner to collect the sandstone samples that would complete their collections.

All crew members agree that each facet of our experience at MDRS contributed to a successful mission. In addition to completing our science goals, we were able to complete all of our Sim goals as well as our collaboration/networking and teambuilding goals. We are grateful for the opportunity to learn and grow in this once-in-a-lifetime experience. Our lives are changed for the better, and we are all eager to return to our classrooms to share our week of Sim on Mars. In addition, we will continue working with Dr Rupert to develop curriculum and materials that will allow MDRS to reach teachers far beyond the scope of the Spaceward Bound Utah program.

 

Operations Report – May 6th

Crew 235 Operations Report 6-MAY-2021

SOL: 4

Name of person filing report: Atila Meszaros

Non-nominal systems: ATV-2

Notes on non-nominal systems: ATV-2 gas pedal system is broken. Once you press it, it would get stuck, instead of retracting, and it could lead to an accident. There’s also a shift problem, but we suspect is related to the gas pedal system interfering with the shift change. As we leave on Saturday, the ATVs will have to be tuned up before next season.

Generator: Working nominally.

Hours run: 0 hours

From what time last night: the generator didn’t run

To what time this morning: the generator didn’t run

List any additional daytime hours when the generator was run: N/A

Solar— SOC 92% at 20:09 (Before generator is run at night)

Notes on power system: Nothing to report

Diesel Reading – 47%

Station Propane Reading – 80%

Outpost Propane Reading – 1. 70% 2. 80%

Water (loft tank): 25 gallons

Water Meter: 151161.3 units

Water (static tank): 250 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Opportunity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: Yes

Hours: 211.9

Beginning charge: 100%

Ending charge: 31%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Yes

Reason for use: EVA#4 and EVA#5

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 2 hours

Notes on ATVs: ATV 2 gs pedal out of service

HabCar used and why, where?: Yes, to go to town for supplies.

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Working nominally.

Summary of suits and radios: All radios nominal, all functional suits currently charging.

Summary of Hab operations: Mouse found in the morning and released during EVA #5.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: General cleaning and bacterial disposal.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Battery replaced and now rotation system is working.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Thank you for being with us tonight, Bernard!

EVA Report – May 6th

Crew 235 EVA Report 06-May-2021

EVA # 6, 7

Author: Jen Carver-Hunter

Purpose of EVA: Continuation of rock sample collection

EVA 6 Start time: 0800
EVA 6 End time: 0900
EVA 7 Start Time: 0930
EVA 7 End time: 1030

Narrative: One of our crew’s goals is to collect rock samples that replicate the rocks found on Mars. For today’s EVA, our targeted specimens were gypsum and sandstone. The first planned stop for the EVA was the overlook at Copernicus Hwy, and the second planned stop on return to the Hab was Cowboy Corner.

We also decided to use these EVAs as an opportunity for the new crew to conduct an EVA without an MDRS-experienced leader. For this reason, we included a map lesson refresher prior to the EVA to ensure that all crew members would be familiar with the geologic features.

The first EVA crew (EVA 6) left the Hab in vehicles at 0800 and followed Cow Dung Rd north to Brahe Hwy. They missed the turn to Brahe Hwy and found themselves at the quarry, where they turned around and made their way to Brahe Hwy.

Then they followed Brahe Hwy to Copernicus Hwy where they stopped for a chance for a 360-degree look at the formations in the area. Then they reboarded vehicles and returned east on Brahe Hwy toward Cow Dung Road until they spotted the glint of the sun shining off the gypsum on the ground.

They collected samples of gypsum near Beige Moon before continuing on to Cowboy Corner, where they decided to not stop for sandstone collection. They also needed to release a packrat near the reservoir. After releasing it, they returned to vehicles and came back to the Hab without incident.

The second EVA crew (EVA 7) left the Hab in vehicles at 0930 and also followed Cow Dung Road north to Brahe Hwy. They did not miss the turn, so they executed a perfect left turn toward Copernicus Hwy, and stopped at the intersection of the two roads to take in the sights at the overlook.

The crew then started up their vehicles (one rover and one ATV) to start their trip back to the Hab. They also stopped near Beige Moon to collect gypsum samples. They elected to also stop at Cowboy Corner to collect sandstone samples before returning to the Hab without incident.

Destination: The overlook at Copernicus Hwy/Brahe Hwy; Beige Moon, Cowboy Corner

Coordinates (use UTM NAD27 CONUS): Destination 1 – 516100, 4254400; Destination 2 – 516500, 4254400; Destination 3 – 518350, 4252700

EVA 4 Participants: Jeff, Randall, Thomas

EVA 5 Participants: Krysta, Jennifer, Allison

Road(s) and routes per MDRS Map: Entrance Rd to Cow Dung Rd to Brahe Hwy

Mode of travel: Pathfinder Rover and 2 ATVs (only 1 ATV used on the second EVA)

Journalist Report – May 6th

Crew 235 Journalist Report 06May2021
Author: Jeff Streba, Crew Journalist
Sol 4: Log Entry

Another beautiful day on Mars.
We had an early morning EVA out to the overlook and then we collected Gypsum and Sandstone samples. Because of the lack of functional suits, only 3 people went on
each EVA.

While one EVA team was out the other team analyzed bacteria. We were all impressed with the results.

When everyone was together, we discussed how to incorporate both the bacteria experiment and rockhounding into our classrooms.

We then ended our sim

Journalist Report – May 5th

Crew 235 Journalist Report 05May2021
Authors: Thomas Quayle and Randall Gibson, Crew Journalist

Sol 3: Generally Speaking, It Goes Like This

Morning began with a brisk wind and another species of Rodentia captured. The crew has been remarkably passive about the recurring nature of the event. What began as a mixed curiosity has evolved into a morning routine, complete with EVA procedures, and smiles.

However, the first order of the day: coffee and another round of lively discussion surrounded the breakfast table. Today’s topic was the politics of Earth. We are hopeful they will come to solve their problems in time. The morning fair consisted of sausage and eggs. It has been great to enjoy the variety of meals that each of the crew members has been willing to jump in and prepare.

Little did the crew suspect that the quiet morning would not remain so for long. Even the most well-planned EVA cannot account for everything a crew can prepare for or expect. Their mission was to search for specific geologic samples in a region that was located nearly twice as far as the previous day’s activities.

More than an hour had passed from when the first EVA team had departed when a call came over the radio, “Hab, Hab!..”. The other half of the crew at the Hab quickly responded and was informed that the EVA was being aborted. Within minutes of the radio communique, the EVA team had arrived.

Thankfully no one had been injured, but a malfunction in the power pack of the Commander’s spacesuit (redacted) had occurred that seriously affected the ability for air to be recirculated. While it is not yet known where the malfunction took place or how it was missed, the event has impacted our EVA assessments for the duration of our mission. Each team will now have to continue with one crew member volunteering to abstain from participation. In the end, the second EVA team was able to successfully manage the completion of objectives scheduled for the day.

After the second EVA returned to the Hab, the crew reconvened as a single group to rest and clean the remains of lunch until the next event of the day. Shannon led a crew discussion on the assembly of Estes Rockets and the pitfalls to consider when introducing the topic to students. Each member of the crew was able to construct their own prototype, which should be available for launch tomorrow.

The crew then joined Shannon in the Science Dome to compare the samples that had been collected across multiple EVA’s. In the last significant event of the day, the astronomy box was retrieved from the outpost. This was the final component required to activate the solar observatory, which the crew is looking forward to using tomorrow.

As a social component, the crew finalized a Lego build that had been started the previous day. The genders and position of the Lego figures were debated and rearranged by the crew. The dinner discussion covered several topics, including the similarities to the Lewis and Clark expedition.

Jennifer prepared an incredible culinary masterpiece of southern fried chicken, fried potatoes, biscuits, and gravy by using dehydrated ingredients. This was surely the best meal yet to be prepared on this planet, and she shall be henceforth be known as the “Mars”tha Stewart.

Sol Summary – May 5th

Crew 235 Sol Summary Report 05-May-2021

Sol: 3

Summary Title: Rocks and Rockets!

Author’s name: Atila Meszaros

Mission Status: Crew Biologist

Sol Activity Summary:

We continued or geology hunting as EVA team #1 and #2 went South and reached the 1101 road to collect granite, lava in its different color variation, asphalt, devil’s toenail, and more to be analyzed by our expert, Dr. Shannon Rupert.

The first EVA encountered a small incident (check the EVA report for full explanation), but we are very close to finishing our geology hunt and gathering.

After lunch, we decided to start building our rockets and launching system for tomorrow’s project! Jennifer and Krysta are making dinner as we prepare the reports and get ready for a great night trying new activities and games.

Look Ahead Plan: Rock hunting EVAs in the morning, rocket launching, and going out of sim!

Anomalies in work: None.

Weather: Fully sunny, but with a great temperature throughout the day.

Crew Physical Status: Healthy. Finally learned how to work with dehydrated food and happy with the great meals and bread.

EVA: Tomorrow EVAs #5 and #6

Reports to be filed: Operations Report, EVA Report, EVA Request, Crew Journalist

Support Requested: Nothing to report.

Operations Report – May 5th

Crew 235 Operations Report 5-MAY-2021

SOL: 3

Name of person filing report: Atila Meszaros

Non-nominal systems: 1 EVA suit, Science Dome door, and inner airlock door of the RAM

Notes on non-nominal systems: Shannon used a screwdriver, a drill, a socket set, and a sledgehammer to fix both doors!

Generator: Working nominally.

Hours run: 0 hours

From what time last night: the generator didn’t run

To what time this morning: the generator didn’t run

List any additional daytime hours when the generator was run: N/A

Solar— SOC 98% at 18:23 (Before generator is run at night)

Notes on power system: Nothing to report

Diesel Reading – 47%

Station Propane Reading – 80%

Outpost Propane Reading – 1. 71% 2. 80%

Water (loft tank): 50 gallons

Water Meter: 150994.0 units

Water (static tank): 350 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Opportunity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: Yes

Hours: 209.2

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Yes

Reason for use: EVA#3 and EVA#4

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 0

Notes on ATVs: ATVs will be filled with ethanol-free gas to be ready for tomorrow’s EVA.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Working nominally.

Summary of suits and radios: All radios nominal, all functional suits currently charging.

Summary of Hab operations: Mouse found in the morning and released during EVA #3.

Summary of GreenHab operations: Door not quite fixed.

Summary of ScienceDome operations: Door fixed.

Summary of RAM operations: Door fixed.

Summary of any observatory issues: The rotation problem was a battery issue.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Thank you for being with us tonight!

EVA Report – May 5th

Crew 235 EVA Report 05 May 2021

EVA # 4, 5

Author: Jen Carver-Hunter

Purpose of EVA: Continuation of rock sample collection

EVA 4 Start time: 0900

EVA 4 End time: 1000

EVA 5 Start Time: 1200

EVA 5 End time: 1430

Narrative: One of our crew’s goals is to collect rock samples that replicate the rocks found on Mars. For today’s EVA, our targeted specimens were lava, basalt, and granite. In addition, we developed a scavenger hunt challenge for the two EVA crews to use as a competition that included specimens of gryphaea, purple lava, hematite, and sandstone blueberries.

The first planned stop for the EVA was the route to Barainca Butte, and the second planned stop on return to the Hab was east of Zubrin’s head.

The first EVA crew left the hab in vehicles at 0900 and followed Cow Dung Road south to Road 1101. After following Road 1101 west toward Barainca Butte, the crew stopped to collect rock samples, which included lava, granite, and basalt, and gryphaea. The Commander’s suit battery discharged so the EVA was aborted and the crew returned to the hab at 1000.

Because that crew’s EVA mission was not completed they requested that the second EVA crew collect extra samples to meet the day’s goals. They also requested that the second crew return some unwanted petrified wood samples to the west side of Kissing Camel Ridge, as was suggested by Dr. Rupert when she approved today’s EVA last night. Due to the dead battery in Spacesuit 7, the second EVA was only able to accommodate 3 crew members. Krysta volunteered to stay at the hab and requested that her crewmates bring samples back for her collection.

The second EVA crew left the hab in vehicles at 1200 and also followed Cow Dung Road south to Road 1101. They followed Road 1101 toward Barainca Butte and collected their samples along the road. After collecting on Road 1101, the EVA team turned around and turned north on Cow Dung Road until they reached “the Squirrel” landmark. They followed a wash to the west (they were east of Zubrin’s head) until they found the sandstone blueberry nursery.

At this point, they were able to complete the second objective of their mission. The final objective was to return the pieces of petrified wood they were done using to their collection site. Crew stopped south of Kissing Camels Ridge and walked west to return their samples before returning back to the Hab. The second EVA crew returned to the Hab with no difficulty and shared their collected samples with the first EVA crew.

Destination: Barainca Butte, east of Zubrin’s head, Kissing Camel Ridge

Coordinates (use UTM NAD27 CONUS): Destination 1 – 519700, 4247200; Destination 2 – 519750, 4248000; Destination 3 – 518250, 4249250

EVA 4 Participants: Jen, Krysta, Jennifer, Jeff

EVA 5 Participants: Atila, Randall, Thomas

Road(s) and routes per MDRS Map: Entrance Road to Cow Dung Road to Road 1101

Mode of travel: Pathfinder Rover and 2 ATVs.